Workers: Akutsu, Michimura, K. Tanaka (10:30-21:00)
GPT part installation
- Other things
- LVDT driver was not working properly last night because the power cable for SG generating 10kHz calibration line for LVDT was accidentaly unplugged. We taped the cable to the reel.
- Height of the BRT output beams to TMSX table was re-adjusted with HBS and GST0, using a laser leveler, so that output beams are parallel to the ground.
- Plugged Dsub cables to green QPDs, and roughly centered (with ceiling LED lights on). Green beam moves a lot.
- I was amazed by how well the cabling was done! We could also see the singal with MEDM screens without any problem.
- The Dsub cables go under the optical table to QPDs.
- Aligned beam to green DC PD to maximize the signal (~12 counts (~6 counts background with ceiling LED lights on)).
- Re-did the cabling of BNC cable around the rack
- Kokeyama-san and T. Yamamoto-san fixed the network issue and Gig-E camera now works. We renamed the camera from BRTX_G(and BRTX_IR) to TMSX_G(and TMSX_IR). The green beam looks clipping somewhere (output of BRT is already dirty).
- The height of Gig-E camera is also adjusted
- Fixed MEDM screen for camera name change, and added a button for GR_PD.
- Installed IR periscope.
- Detach the remaining white tapes just before closing the chamber; be noted that there are white tapes on the PEEK connectors at the inside of the flanges.
- Torque completely the bolts which are fixing the upper and the lower base plates.
- Insert metal gaskets before torquing the ICF203 flanges with Dsub connectors.
- Attach an oplev viewport
- Setup the oplev for TMS-VIS
- Consider why the green ghost beam is not observed; that was observed in TMSY in this April.
- Modify MEDM screen
- Add DC PD and so on.
Cable Dsub 15pin for TMS-VIS OpLev.
- Return plastic boxes.
- Install rails in digital rack.
- They may be necessary under LVDT distributor, geophone distributor, and WHT board.
- Nagano is not sure it is necessary under HP coil driver.
- Consider how to cable Dsub cables to QPD.
- Install IR optics
- Measure the length between BRT lens3 and GLNS1
- Black tubes attachable to the DC-PDs or seetabale in front of the DC-QPDs for light shielding.
- Screws to screw Dsub 15pin connectors to QPD amplifiers
- Nice way to turn off the ceiling LED light
- Covers for cables running on the floor
- Shield box against environmental light
- Wind shielding... (it would be somewhat difficult as two beams are running in the different directions from the viewport window)